\begin{thebibliography}{}

\bibitem[\protect\citename{Angeli and
  Manning}2013]{angeli-manning:2013:CoNLL-2013}
Gabor Angeli and Christopher Manning.
\newblock 2013.
\newblock {Philosophers are Mortal: Inferring the Truth of Unseen Facts}.
\newblock In {\em Proceedings of the Seventeenth Conference on Computational
  Natural Language Learning}, pages 133--142.

\bibitem[\protect\citename{Bollacker \bgroup et al.\egroup
  }2008]{Bollacker:2008}
Kurt Bollacker, Colin Evans, Praveen Paritosh, Tim Sturge, and Jamie Taylor.
\newblock 2008.
\newblock {Freebase: A Collaboratively Created Graph Database for Structuring
  Human Knowledge}.
\newblock In {\em Proceedings of the 2008 ACM SIGMOD International Conference
  on Management of Data}, pages 1247--1250.

\bibitem[\protect\citename{Bordes \bgroup et al.\egroup }2011]{bordes-2011}
Antoine Bordes, Jason Weston, Ronan Collobert, and Yoshua Bengio.
\newblock 2011.
\newblock {Learning Structured Embeddings of Knowledge Bases}.
\newblock In {\em Proceedings of the Twenty-Fifth AAAI Conference on Artificial
  Intelligence}, pages 301--306.

\bibitem[\protect\citename{Bordes \bgroup et al.\egroup }2012]{Bordes2014SME}
Antoine Bordes, Xavier Glorot, Jason Weston, and Yoshua Bengio.
\newblock 2012.
\newblock {A Semantic Matching Energy Function for Learning with
  Multi-relational Data}.
\newblock {\em Machine Learning}, 94(2):233--259.

\bibitem[\protect\citename{Bordes \bgroup et al.\egroup }2013]{NIPS2013_5071}
Antoine Bordes, Nicolas Usunier, Alberto Garcia-Duran, Jason Weston, and Oksana
  Yakhnenko.
\newblock 2013.
\newblock {Translating Embeddings for Modeling Multi-relational Data}.
\newblock In {\em Advances in Neural Information Processing Systems 26}, pages
  2787--2795.

\bibitem[\protect\citename{Das \bgroup et al.\egroup }2017]{DasNBM17}
Rajarshi Das, Arvind Neelakantan, David Belanger, and Andrew McCallum.
\newblock 2017.
\newblock Chains of reasoning over entities, relations, and text using
  recurrent neural networks.
\newblock In {\em Proceedings of the 15th Conference of the European Chapter of
  the Association for Computational Linguistics}.

\bibitem[\protect\citename{Duchi \bgroup et al.\egroup
  }2011]{Duchi:2011:ASM:1953048.2021068}
John Duchi, Elad Hazan, and Yoram Singer.
\newblock 2011.
\newblock {Adaptive Subgradient Methods for Online Learning and Stochastic
  Optimization}.
\newblock {\em The Journal of Machine Learning Research}, 12:2121--2159.

\bibitem[\protect\citename{Fellbaum}1998]{FellbaumC98}
Christiane~D. Fellbaum.
\newblock 1998.
\newblock {\em {WordNet}: An Electronic Lexical Database}.
\newblock MIT Press.

\bibitem[\protect\citename{Feng \bgroup et al.\egroup
  }2016]{feng-EtAl:2016:COLING1}
Jun Feng, Minlie Huang, Yang Yang, and xiaoyan zhu.
\newblock 2016.
\newblock {GAKE: Graph Aware Knowledge Embedding}.
\newblock In {\em Proceedings of COLING 2016, the 26th International Conference
  on Computational Linguistics: Technical Papers}, pages 641--651.

\bibitem[\protect\citename{Garc{\'{\i}}a{-}Dur{\'{a}}n \bgroup et al.\egroup
  }2015]{garciaduran-bordes-usunier:2015:EMNLP}
Alberto Garc{\'{\i}}a{-}Dur{\'{a}}n, Antoine Bordes, and Nicolas Usunier.
\newblock 2015.
\newblock {Composing Relationships with Translations}.
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 286--290.

\bibitem[\protect\citename{Garc{\'{\i}}a{-}Dur{\'{a}}n \bgroup et al.\egroup
  }2016]{Garcia-DuranBUG15}
Alberto Garc{\'{\i}}a{-}Dur{\'{a}}n, Antoine Bordes, Nicolas Usunier, and Yves
  Grandvalet.
\newblock 2016.
\newblock {Combining Two and Three-Way Embedding Models for Link Prediction in
  Knowledge Bases}.
\newblock {\em Journal of Artificial Intelligence Research}, 55:715--742.

\bibitem[\protect\citename{Guu \bgroup et al.\egroup
  }2015]{guu-miller-liang:2015:EMNLP}
Kelvin Guu, John Miller, and Percy Liang.
\newblock 2015.
\newblock {Traversing Knowledge Graphs in Vector Space}.
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 318--327.

\bibitem[\protect\citename{He \bgroup et al.\egroup }2015]{He:2015}
Shizhu He, Kang Liu, Guoliang Ji, and Jun Zhao.
\newblock 2015.
\newblock {Learning to Represent Knowledge Graphs with Gaussian Embedding}.
\newblock In {\em Proceedings of the 24th ACM International on Conference on
  Information and Knowledge Management}, pages 623--632.

\bibitem[\protect\citename{Jenatton \bgroup et al.\egroup }2012]{NIPS2012_4744}
Rodolphe Jenatton, Nicolas~L. Roux, Antoine Bordes, and Guillaume~R Obozinski.
\newblock 2012.
\newblock {A latent factor model for highly multi-relational data}.
\newblock In {\em Advances in Neural Information Processing Systems 25}, pages
  3167--3175.

\bibitem[\protect\citename{Ji \bgroup et al.\egroup
  }2015]{ji-EtAl:2015:ACL-IJCNLP}
Guoliang Ji, Shizhu He, Liheng Xu, Kang Liu, and Jun Zhao.
\newblock 2015.
\newblock {Knowledge Graph Embedding via Dynamic Mapping Matrix}.
\newblock In {\em Proceedings of the 53rd Annual Meeting of the Association for
  Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 687--696.

\bibitem[\protect\citename{Ji \bgroup et al.\egroup }2016]{JiLH016}
Guoliang Ji, Kang Liu, Shizhu He, and Jun Zhao.
\newblock 2016.
\newblock {Knowledge Graph Completion with Adaptive Sparse Transfer Matrix}.
\newblock In {\em Proceedings of the Thirtieth {AAAI} Conference on Artificial
  Intelligence}, pages 985--991.

\bibitem[\protect\citename{Krompa{\ss} \bgroup et al.\egroup }2015]{Denis2015}
Denis Krompa{\ss}, Stephan Baier, and Volker Tresp.
\newblock 2015.
\newblock {Type-Constrained Representation Learning in Knowledge Graphs}.
\newblock In {\em Proceedings of the 14th International Semantic Web
  Conference}, pages 640--655.

\bibitem[\protect\citename{Lehmann \bgroup et al.\egroup
  }2015]{LehmannIJJKMHMK15}
Jens Lehmann, Robert Isele, Max Jakob, Anja Jentzsch, Dimitris Kontokostas,
  Pablo~N. Mendes, Sebastian Hellmann, Mohamed Morsey, Patrick van Kleef,
  S{\"{o}}ren Auer, and Christian Bizer.
\newblock 2015.
\newblock {DBpedia - A Large-scale, Multilingual Knowledge Base Extracted from
  Wikipedia}.
\newblock {\em Semantic Web}, 6(2):167--195.

\bibitem[\protect\citename{Lin \bgroup et al.\egroup
  }2015a]{lin-EtAl:2015:EMNLP1}
Yankai Lin, Zhiyuan Liu, Huanbo Luan, Maosong Sun, Siwei Rao, and Song Liu.
\newblock 2015a.
\newblock {Modeling Relation Paths for Representation Learning of Knowledge
  Bases}.
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 705--714.

\bibitem[\protect\citename{Lin \bgroup et al.\egroup }2015b]{AAAI159571}
Yankai Lin, Zhiyuan Liu, Maosong Sun, Yang Liu, and Xuan Zhu.
\newblock 2015b.
\newblock {Learning Entity and Relation Embeddings for Knowledge Graph
  Completion}.
\newblock In {\em Proceedings of the Twenty-Ninth AAAI Conference on Artificial
  Intelligence Learning}, pages 2181--2187.

\bibitem[\protect\citename{Liu and Nocedal}1989]{Liu1989}
D.~C. Liu and J.~Nocedal.
\newblock 1989.
\newblock {On the Limited Memory BFGS Method for Large Scale Optimization}.
\newblock {\em Mathematical Programming}, 45(3):503--528.

\bibitem[\protect\citename{Liu \bgroup et al.\egroup
  }2016]{Liu:2016:HRW:2911451.2911509}
Qiao Liu, Liuyi Jiang, Minghao Han, Yao Liu, and Zhiguang Qin.
\newblock 2016.
\newblock {Hierarchical Random Walk Inference in Knowledge Graphs}.
\newblock In {\em Proceedings of the 39th International ACM SIGIR Conference on
  Research and Development in Information Retrieval}, pages 445--454.

\bibitem[\protect\citename{Mikolov \bgroup et al.\egroup }2013]{Mikolov13b}
Tomas Mikolov, Wen-tau Yih, and Geoffrey Zweig.
\newblock 2013.
\newblock {Linguistic Regularities in Continuous Space Word Representations}.
\newblock In {\em Proceedings of the 2013 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 746--751.

\bibitem[\protect\citename{Nguyen \bgroup et al.\egroup }2016]{NguyenCoNLL2016}
Dat~Quoc Nguyen, Kairit Sirts, Lizhen Qu, and Mark Johnson.
\newblock 2016.
\newblock {Neighborhood Mixture Model for Knowledge Base Completion}.
\newblock In {\em Proceedings of The 20th SIGNLL Conference on Computational
  Natural Language Learning}, pages 40--50.

\bibitem[\protect\citename{Nickel \bgroup et al.\egroup
  }2011]{ICML2011Nickel_438}
Maximilian Nickel, Volker Tresp, and Hans-Peter Kriegel.
\newblock 2011.
\newblock {A Three-Way Model for Collective Learning on Multi-Relational Data}.
\newblock In {\em Proceedings of the 28th International Conference on Machine
  Learning}, pages 809--816.

\bibitem[\protect\citename{Nickel \bgroup et al.\egroup }2016a]{NickelMTG15}
Maximilian Nickel, Kevin Murphy, Volker Tresp, and Evgeniy Gabrilovich.
\newblock 2016a.
\newblock {A Review of Relational Machine Learning for Knowledge Graphs}.
\newblock {\em Proceedings of the IEEE}, 104(1):11--33.

\bibitem[\protect\citename{Nickel \bgroup et al.\egroup
  }2016b]{Nickel:2016:HEK:3016100.3016172}
Maximilian Nickel, Lorenzo Rosasco, and Tomaso Poggio.
\newblock 2016b.
\newblock Holographic embeddings of knowledge graphs.
\newblock In {\em Proceedings of the Thirtieth AAAI Conference on Artificial
  Intelligence}, pages 1955--1961.

\bibitem[\protect\citename{Niepert}2016]{NIPS2016_6098}
Mathias Niepert.
\newblock 2016.
\newblock {Discriminative Gaifman Models}.
\newblock In {\em Advances in Neural Information Processing Systems 29}, pages
  3405--3413.

\bibitem[\protect\citename{Riedel \bgroup et al.\egroup
  }2013]{riedel-EtAl:2013:NAACL-HLT}
Sebastian Riedel, Limin Yao, Andrew McCallum, and Benjamin~M. Marlin.
\newblock 2013.
\newblock {Relation Extraction with Matrix Factorization and Universal
  Schemas}.
\newblock In {\em Proceedings of the 2013 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 74--84.

\bibitem[\protect\citename{Shi and Weninger}2017]{ShiW16a}
Baoxu Shi and Tim Weninger.
\newblock 2017.
\newblock {ProjE: Embedding Projection for Knowledge Graph Completion}.
\newblock In {\em Proceedings of the 31st AAAI Conference on Artificial
  Intelligence}.

\bibitem[\protect\citename{Socher \bgroup et al.\egroup }2013]{NIPS2013_5028}
Richard Socher, Danqi Chen, Christopher~D Manning, and Andrew Ng.
\newblock 2013.
\newblock {Reasoning With Neural Tensor Networks for Knowledge Base
  Completion}.
\newblock In {\em Advances in Neural Information Processing Systems 26}, pages
  926--934.

\bibitem[\protect\citename{Suchanek \bgroup et al.\egroup }2007]{Suchanek:2007}
Fabian~M. Suchanek, Gjergji Kasneci, and Gerhard Weikum.
\newblock 2007.
\newblock {YAGO: A Core of Semantic Knowledge}.
\newblock In {\em Proceedings of the 16th International Conference on World
  Wide Web}, pages 697--706.

\bibitem[\protect\citename{Taskar \bgroup et al.\egroup }2004]{NIPS2003_2465}
Ben Taskar, Ming fai Wong, Pieter Abbeel, and Daphne Koller.
\newblock 2004.
\newblock {Link Prediction in Relational Data}.
\newblock In {\em Advances in Neural Information Processing Systems 16}, pages
  659--666.

\bibitem[\protect\citename{Tay \bgroup et al.\egroup
  }2017]{Tay:2017:RST:3018661.3018695}
Yi~Tay, Anh~Tuan Luu, Siu~Cheung Hui, and Falk Brauer.
\newblock 2017.
\newblock {Random Semantic Tensor Ensemble for Scalable Knowledge Graph Link
  Prediction}.
\newblock In {\em Proceedings of the Tenth ACM International Conference on Web
  Search and Data Mining}, pages 751--760.

\bibitem[\protect\citename{Toutanova and Chen}2015]{toutanova-chen:2015:CVSC}
Kristina Toutanova and Danqi Chen.
\newblock 2015.
\newblock {Observed Versus Latent Features for Knowledge Base and Text
  Inference}.
\newblock In {\em Proceedings of the 3rd Workshop on Continuous Vector Space
  Models and their Compositionality}, pages 57--66.

\bibitem[\protect\citename{Toutanova \bgroup et al.\egroup
  }2015]{toutanova-EtAl:2015:EMNLP}
Kristina Toutanova, Danqi Chen, Patrick Pantel, Hoifung Poon, Pallavi
  Choudhury, and Michael Gamon.
\newblock 2015.
\newblock {Representing Text for Joint Embedding of Text and Knowledge Bases}.
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 1499--1509.

\bibitem[\protect\citename{Toutanova \bgroup et al.\egroup
  }2016]{toutanova-EtAl:2016:P16-1}
Kristina Toutanova, Victoria Lin, Wen-tau Yih, Hoifung Poon, and Chris Quirk.
\newblock 2016.
\newblock {Compositional Learning of Embeddings for Relation Paths in Knowledge
  Base and Text}.
\newblock In {\em Proceedings of the 54th Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers)}, pages 1434--1444.

\bibitem[\protect\citename{Trouillon \bgroup et al.\egroup
  }2016]{TrouillonWRGB16}
Th{\'{e}}o Trouillon, Johannes Welbl, Sebastian Riedel, {\'{E}}ric Gaussier,
  and Guillaume Bouchard.
\newblock 2016.
\newblock {Complex Embeddings for Simple Link Prediction}.
\newblock In {\em Proceedings of the 33nd International Conference on Machine
  Learning}, pages 2071--2080.

\bibitem[\protect\citename{Wang and Li}2016]{DBLP:conf/ijcai/WangL16}
Zhigang Wang and Juan{-}Zi Li.
\newblock 2016.
\newblock {Text-Enhanced Representation Learning for Knowledge Graph}.
\newblock In {\em Proceedings of the Twenty-Fifth International Joint
  Conference on Artificial Intelligence}, pages 1293--1299.

\bibitem[\protect\citename{Wang \bgroup et al.\egroup
  }2014a]{wang-EtAl:2014:EMNLP20145}
Zhen Wang, Jianwen Zhang, Jianlin Feng, and Zheng Chen.
\newblock 2014a.
\newblock {Knowledge Graph and Text Jointly Embedding}.
\newblock In {\em Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 1591--1601.

\bibitem[\protect\citename{Wang \bgroup et al.\egroup }2014b]{AAAI148531}
Zhen Wang, Jianwen Zhang, Jianlin Feng, and Zheng Chen.
\newblock 2014b.
\newblock {Knowledge Graph Embedding by Translating on Hyperplanes}.
\newblock In {\em Proceedings of the Twenty-Eighth AAAI Conference on
  Artificial Intelligence}, pages 1112--1119.

\bibitem[\protect\citename{Wang \bgroup et al.\egroup
  }2016]{wang-EtAl:2016:P16-13}
Quan Wang, Jing Liu, Yuanfei Luo, Bin Wang, and Chin-Yew Lin.
\newblock 2016.
\newblock {Knowledge Base Completion via Coupled Path Ranking}.
\newblock In {\em Proceedings of the 54th Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers)}, pages 1308--1318.

\bibitem[\protect\citename{Wei \bgroup et al.\egroup
  }2016]{wei-zhao-liu:2016:EMNLP2016}
Zhuoyu Wei, Jun Zhao, and Kang Liu.
\newblock 2016.
\newblock {Mining Inference Formulas by Goal-Directed Random Walks}.
\newblock In {\em Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pages 1379--1388.

\bibitem[\protect\citename{West \bgroup et al.\egroup
  }2014]{West:2014:KBC:2566486.2568032}
Robert West, Evgeniy Gabrilovich, Kevin Murphy, Shaohua Sun, Rahul Gupta, and
  Dekang Lin.
\newblock 2014.
\newblock {Knowledge Base Completion via Search-based Question Answering}.
\newblock In {\em Proceedings of the 23rd International Conference on World
  Wide Web}, pages 515--526.

\bibitem[\protect\citename{Xiao \bgroup et al.\egroup }2017]{0005HZ16}
Han Xiao, Minlie Huang, and Xiaoyan Zhu.
\newblock 2017.
\newblock {SSP:} semantic space projection for knowledge graph embedding with
  text descriptions.
\newblock In {\em Proceedings of the 31st AAAI Conference on Artificial
  Intelligence}.

\bibitem[\protect\citename{Yang \bgroup et al.\egroup }2015]{yang-etal-2015}
Bishan Yang, Wen-tau Yih, Xiaodong He, Jianfeng Gao, and Li~Deng.
\newblock 2015.
\newblock {Embedding Entities and Relations for Learning and Inference in
  Knowledge Bases}.
\newblock In {\em Proceedings of the International Conference on Learning
  Representations}.

\bibitem[\protect\citename{Yoon \bgroup et al.\egroup
  }2016]{yoon-EtAl:2016:N16-1}
Hee-Geun Yoon, Hyun-Je Song, Seong-Bae Park, and Se-Young Park.
\newblock 2016.
\newblock {A Translation-Based Knowledge Graph Embedding Preserving Logical
  Property of Relations}.
\newblock In {\em Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 907--916.

\bibitem[\protect\citename{Zeiler}2012]{DBLP:journals/corr/abs-1212-5701}
Matthew~D. Zeiler.
\newblock 2012.
\newblock {{ADADELTA:} An Adaptive Learning Rate Method}.
\newblock {\em CoRR}, abs/1212.5701.

\bibitem[\protect\citename{Zhao \bgroup et al.\egroup }2015]{zhao2015y}
Yu~Zhao, Sheng Gao, Patrick Gallinari, and Jun Guo.
\newblock 2015.
\newblock {Knowledge Base Completion by Learning Pairwise-Interaction
  Differentiated Embeddings}.
\newblock {\em Data Mining and Knowledge Discovery}, 29(5):1486--1504.

\end{thebibliography}

\documentclass[11pt,letterpaper]{article}
\pdfoutput=1
\usepackage{naaclhlt2016}
\usepackage{times}
\usepackage{latexsym}
\usepackage{amsmath}
\usepackage{multirow}
\usepackage{url}
\usepackage{graphicx}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage[hidelinks]{hyperref}
\usepackage{fancyvrb}
\newcommand{\citep}[1]{\cite{#1}}
\newcommand{\citet}[1]{\newcite{#1}}

\naaclfinalcopy 
\def\naaclpaperid{132} 

\usepackage{xcolor}
\newcommand{\CHANGEA}[1]{#1}   
\newcommand{\CHANGEB}[1]{#1}   

\title{STransE: a novel embedding model of entities and relationships \\ in knowledge bases\thanks{\ \ A revised version of our NAACL-HLT 2016 paper with additional experimental results and latest related work.}}

\author{Dat Quoc Nguyen${}^{1}$, Kairit Sirts${}^{1}$, Lizhen Qu${}^{2}$ \and Mark Johnson${}^{1}$ \\
\\
${}^{1}$ Department of Computing, Macquarie University, Sydney, Australia \\
{\tt{\small{dat.nguyen@students.mq.edu.au, \{kairit.sirts, mark.johnson\}@mq.edu.au}}} \\
${}^{2}$ NICTA, ACT 2601, Australia \\
{\tt{\small{lizhen.qu@nicta.com.au}}}
}

\begin{document}

\maketitle

\begin{abstract}
Knowledge bases of real-world facts about entities and their relationships are useful resources for a variety of natural language processing tasks. However, because knowledge bases are typically incomplete, it is useful to be able to perform \textit{link prediction} or \emph{knowledge base completion}, i.e., predict whether a relationship not in the knowledge base is likely to be true.  \CHANGEA{This paper combines insights \CHANGEB{from} several previous link prediction models into a new embedding model \textit{STransE}} that represents each entity as a low-dimensional vector, and each relation by two matrices and a translation vector. \CHANGEA{STransE is \CHANGEB{a simple combination of the SE and TransE  models,} but} \CHANGEA{it} obtains better link prediction performance on two benchmark datasets than previous embedding models. \CHANGEA{Thus, \CHANGEB{STransE can serve as} a new baseline for the more complex models in \CHANGEB{the} link prediction task.}
\end{abstract}

\section{Introduction} \label{sec:intro}
Knowledge bases (KBs), such as WordNet \cite{FellbaumC98}, YAGO
\cite{Suchanek:2007}, Freebase \cite{Bollacker:2008} and DBpedia
\cite{LehmannIJJKMHMK15}, represent relationships between entities as
triples $(\mathrm{head\ entity, relation, tail\ entity})$.  Even very
large knowledge bases are still far from complete
\cite{NIPS2013_5028,West:2014:KBC:2566486.2568032}.
\emph{Link prediction} or \emph{knowledge base completion} systems
\cite{NickelMTG15} predict which triples not in a knowledge base are
likely to be true  \cite{NIPS2003_2465,bordes-2011}.  A variety of
different kinds of information is potentially useful here, including
information extracted from external corpora
\cite{riedel-EtAl:2013:NAACL-HLT,wang-EtAl:2014:EMNLP20145} and the
other relationships that hold between the entities
\cite{angeli-manning:2013:CoNLL-2013,zhao2015y}.
\CHANGEB{For example, \newcite{toutanova-EtAl:2015:EMNLP} used information
from the external ClueWeb-12 corpus to significantly enhance performance.}

While integrating a wide variety of information sources can produce excellent results \citep{DasNBM17},
there are several reasons for studying simpler models that
directly optimize a score function for the triples in a knowledge base, such as the one presented here.
\CHANGEB{First, additional information sources might not be available, e.g., for knowledge bases for specialized domains.}
\CHANGEB{Second}, models \CHANGEB{that don't exploit external resources} are simpler and thus typically much faster to train than the more complex models using additional information. 
\CHANGEB{Third}, the more complex models \CHANGEB{that exploit external information} 
are typically extensions of these simpler models,
and are often initialized with parameters estimated by such simpler models,
so improvements to the simpler models should yield corresponding improvements
to the more complex models \CHANGEB{as well}.

\begin{table*}[ht]
\centering
{\footnotesize
\resizebox{16.5cm}{!}{
\def\arraystretch{1.25}
\begin{tabular}{l|l|l}
\hline
\bf Model & Score function $f_r(h, t)$ & \bf Opt.  \\
\hline
\hline
SE &  $\| \textbf{W}_{r,1}\textbf{h} - \textbf{W}_{r,2}\textbf{t}\|_{\ell_{1/2}}$ ;  $\textbf{W}_{r,1}$, $\textbf{W}_{r,2}$ $\in$ $\mathbb{R}^{k \times k}$ & SGD\\
\hline
Unstructured &  $ \| \textbf{h} - \textbf{t} \|_{\ell_{1/2}}$ & SGD \\
\hline
TransE &  $ \|  \textbf{h} + \textbf{r} - \textbf{t} \|_{\ell_{1/2}}$ ;    \textbf{r} $\in$ $\mathbb{R}^{k}$ & SGD \\
\hline
DISTMULT &  $ \textbf{h}^{\top}\textbf{W}_{r}\textbf{t}$ ;   $\textbf{W}_{r}$ is a diagonal matrix $\in$ $\mathbb{R}^{k \times k}$ & AdaGrad\\
\hline
{NTN} & $ \textbf{u}_r^{\top} \mathit{tanh}( \textbf{h}^{\top}\textbf{M}_{r}\textbf{t}  + \textbf{W}_{r,1}\textbf{h} + \textbf{W}_{r,2}\textbf{t} + \textbf{b}_r)$ ; $\textbf{u}_r\text{, } \textbf{b}_r \in \mathbb{R}^d$; $\textbf{M}_{r} \in \mathbb{R}^{k \times k \times d}$;  $\textbf{W}_{r,1}$, $\textbf{W}_{r,2} \in \mathbb{R}^{d \times k}$ & {L-BFGS}  \\
\hline
{TransH} & $\| (\textbf{I} - \textbf{r}_p\textbf{r}_p^{\top})\textbf{h} + \textbf{r} - (\textbf{I} - \textbf{r}_p\textbf{r}_p^{\top})\textbf{t} \|_{\ell_{1/2}}$ ; $\textbf{r}_p$, $\textbf{r} \in$ $\mathbb{R}^{k}$ ; $\textbf{I}$: Identity matrix size $k \times k$ & {SGD}\\
\hline
{TransD} & $\| (\textbf{I} + \textbf{r}_p\textbf{h}_p^{\top})\textbf{h} + \textbf{r} - (\textbf{I} + \textbf{r}_p\textbf{t}_p^{\top})\textbf{t} \|_{\ell_{1/2}}$ ;  \textbf{r}$_p$, \textbf{r} $\in$ $\mathbb{R}^{d}$ ; \textbf{h}$_p$, \textbf{t}$_p$ $\in$ $\mathbb{R}^{k}$ ; $\textbf{I}$: Identity matrix size $d \times k$  & {AdaDelta} \\
\hline
TransR & $\| \textbf{W}_{r}\textbf{h} + \textbf{r} - \textbf{W}_{r}\textbf{t}\|_{\ell_{1/2}}$ ;  $\textbf{W}_{r}$ $\in$ $\mathbb{R}^{d \times k}$  ;    \textbf{r} $\in$ $\mathbb{R}^{d}$ & SGD \\
\hline
TranSparse & $\| \textbf{W}_{r}^h(\theta_r^h)\textbf{h} + \textbf{r} - \textbf{W}_{r}^t(\theta_r^t)\textbf{t} \|_{\ell_{1/2}}$ ;  $\textbf{W}_{r}^h$, $\textbf{W}_{r}^t$ $\in$ $\mathbb{R}^{d \times k}$; $\theta_r^h$, $\theta_r^t \in \mathbb{R}$ ; $\textbf{r} \in$ $\mathbb{R}^{d}$ & SGD  \\
\hline
\hline
Our STransE & $\| \textbf{W}_{r,1}\textbf{h} + \textbf{r} - \textbf{W}_{r,2}\textbf{t}\|_{\ell_{1/2}}$ ;  $\textbf{W}_{r,1}$, $\textbf{W}_{r,2}$ $\in$ $\mathbb{R}^{k \times k}$; \textbf{r} $\in$ $\mathbb{R}^{k}$ & SGD  \\
\hline
\end{tabular}
}
}
\caption{The score functions $f_r(h, t)$ and the optimization methods
  (Opt.) of several prominent embedding models for KB completion.  
  In all of these the entities $h$ and $t$ are represented
  by  vectors $\textbf{h}$ and $\textbf{t} \in \mathbb{R}^{k}$ respectively.
}
\label{tab:emmethods}
\end{table*}

\emph{Embedding models} for KB completion associate entities and/or
relations with dense feature vectors or matrices.  Such models obtain
state-of-the-art performance
\cite{ICML2011Nickel_438,bordes-2011,Bordes2014SME,NIPS2013_5071,NIPS2013_5028,AAAI148531,guu-miller-liang:2015:EMNLP}
and generalize to large KBs \cite{Denis2015}.
Table~\ref{tab:emmethods} \CHANGEB{summarizes} a number of prominent embedding
models for KB completion.

Let $(h, r, t)$ represent a triple.  In all of the models discussed
here, the head entity $h$ and the tail entity $t$ are represented by
vectors $\textbf{h}$ and $\textbf{t}\in\mathbb{R}^{k}$ respectively.
The \emph{Unstructured} model \cite{Bordes2014SME}
assumes that $\textbf{h} \approx \textbf{t}$. As the Unstructured
model does not take the relationship $r$ into account, it cannot
distinguish different relation types.  The \emph{Structured
  Embedding} (SE) model \cite{bordes-2011} extends the unstructured
model by assuming that $h$ and $t$ are similar only in a
relation-dependent subspace.  It represents each relation $r$ with two
matrices $\textbf{W}_{r,1}$ and $\textbf{W}_{r,2}\in\mathbb{R}^{k\times k}$, which are chosen
so that $\textbf{W}_{r,1}\textbf{h} \approx \textbf{W}_{r,2}\textbf{t}$. The \emph{TransE} model
\cite{NIPS2013_5071} is inspired by models such as Word2Vec
\cite{Mikolov13b} where relationships between words often correspond
to translations in latent feature space.  The TransE model
represents each relation $r$ by a translation
vector \textbf{r} $\in\mathbb{R}^{k}$, which is chosen so that 
$\textbf{h} + \textbf{r} \approx \textbf{t}$.

 The primary contribution of \CHANGEB{this} paper is that two very simple relation-prediction models, SE and TransE, can be combined into a single model, which we call \emph{STransE}.\footnote{Source code: \url{https://github.com/datquocnguyen/STransE}}  
Specifically, we use relation-specific matrices $\textbf{W}_{r,1}$ and
$\textbf{W}_{r,2}$ as in the SE model to identify the
relation-dependent aspects of both $h$ and $t$, and use a vector
$\textbf{r}$ as in the TransE model to describe the relationship
between $h$ and $t$ in this subspace.  Specifically, our new KB completion model STransE 
chooses $\textbf{W}_{r,1}$, $\textbf{W}_{r,2}$ and $\textbf{r}$ so that
$\textbf{W}_{r,1}\textbf{h} + \textbf{r} \approx
\textbf{W}_{r,2}\textbf{t}$\CHANGEB{.  That is, a TransE-style relationship} \CHANGEA{holds in some relation-dependent subspace, and crucially, this subspace may involve very different projections of the head $h$ and tail $t$. So $\textbf{W}_{r,1}$ and $\textbf{W}_{r,2}$ can highlight, suppress, or even change the sign of, relation-specific attributes of $h$ and $t$.  For example, for the ``purchases'' relationship, certain attributes of individuals $h$ (e.g., age, gender, marital status) are presumably strongly correlated with very different attributes of objects $t$ (e.g., sports car, washing machine and the like).}

 As we show below,  STransE 
performs better than the SE and TransE models and other
state-of-the-art link prediction models on two standard link prediction datasets WN18 and FB15k, \CHANGEA{so it can serve as a new baseline for KB completion.  We \CHANGEB{expect} that the STransE \CHANGEB{will also be able to} serve as the basis for extended models that exploit a wider variety of information sources\CHANGEB{,} just as TransE does.}

\section{Our approach}
Let $\mathcal{E}$ denote the set of entities and $\mathcal{R}$ the set
of relation types. For each triple $(h, r, t)$, where $h, t \in
\mathcal{E}$ and $r \in \mathcal{R}$, the STransE model defines a \emph{score
function} $f_r(h, t)$ of its implausibility.
Our goal is to choose $f$ such that the score $f_r(h,t)$ of a plausible triple $(h,r,t)$ is
smaller than the score $f_{r'}(h',t')$ of an implausible triple $(h',r',t')$. We define the STransE score function $f$ as follows:

%
\vspace{-15pt}

\begin{eqnarray*}
f_r(h, t) & = & \| \textbf{W}_{r,1}\textbf{h} + \textbf{r} - \textbf{W}_{r,2}\textbf{t}\|_{\ell_{1/2}} 
\end{eqnarray*}


\noindent using either  the $\ell_1$ or the $\ell_2$-norm \CHANGEB{(the choice is made using validation data; in our experiments we found that the $\ell_1$ norm gave slightly better results)}. 
%
 To learn the vectors and matrices we minimize the following margin-based objective function:
%


\begin{eqnarray*}
\mathcal{L} & = & \sum_{\substack{(h,r,t) \in \mathcal{G} \\ (h',r,t') \in \mathcal{G}'_{(h, r, t)}}} [\gamma + f_r(h, t) - f_r(h', t')]_+
\end{eqnarray*}
%



\noindent where $[x]_+ = \max(0, x)$, $\gamma$ is the margin hyper-parameter, $\mathcal{G}$ is
the training set consisting of correct triples, and
$\mathcal{G}'_{(h, r, t)} = \lbrace (h', r, t) \mid h' \in
\mathcal{E}, (h', r, t) \notin \mathcal{G} \rbrace \cup \lbrace (h, r,
t') \mid t' \in \mathcal{E}, (h, r, t') \notin \mathcal{G} \rbrace $
is the set of incorrect triples generated by corrupting a
correct triple $(h, r, t)\in\mathcal{G}$.  

We use Stochastic
Gradient Descent (SGD) to minimize $\mathcal{L}$, and
impose the following
constraints during training:
$\|\textbf{h}\|_2 \leqslant 1$, $\|\textbf{r}\|_2 \leqslant 1$,
$\|\textbf{t}\|_2 \leqslant 1$, $\| \textbf{W}_{r,1}\textbf{h}\|_2
\leqslant 1$ and $\| \textbf{W}_{r,2}\textbf{t}\|_2 \leqslant 1$.
%
%

\section{Related work}
\label{sec:related}
Table \ref{tab:emmethods} summarizes related embedding models for link
prediction and KB completion.  The models differ in the score
functions $f_r(h, t)$ and the algorithms used to optimize the
margin-based objective function, e.g., SGD, AdaGrad
\cite{Duchi:2011:ASM:1953048.2021068}, AdaDelta
\cite{DBLP:journals/corr/abs-1212-5701} and L-BFGS \cite{Liu1989}.

DISTMULT \cite{yang-etal-2015} is based on a Bilinear model
\cite{ICML2011Nickel_438,Bordes2014SME,NIPS2012_4744}
where each relation is represented by a diagonal rather than a full
matrix. 
The neural tensor network (NTN) model \cite{NIPS2013_5028} uses a
bilinear tensor operator to represent each relation while ProjE \citep{ShiW16a} could be viewed as a simplified version of NTN with diagonal matrices. Similar
quadratic forms are used to model entities and relations in  KG2E \citep{He:2015}, ComplEx \citep{TrouillonWRGB16}, TATEC \citep{Garcia-DuranBUG15} and RSTE \citep{Tay:2017:RST:3018661.3018695}. In addition,  HolE \citep{Nickel:2016:HEK:3016100.3016172} uses  circular correlation---a compositional operator---which could be interpreted as a compression of the tensor product. 

The TransH model \cite{AAAI148531} associates each relation with a
relation-specific hyperplane and uses a projection vector to project
entity vectors onto that hyperplane. TransD
\cite{ji-EtAl:2015:ACL-IJCNLP} and TransR/CTransR \cite{AAAI159571}
extend the TransH model using two projection vectors and a matrix to project entity vectors into
a relation-specific space, respectively. TransD learns a relation-role
specific mapping just as STransE, but represents this mapping by projection
vectors rather than full matrices, as in STransE. The lppTransD  model \citep{yoon-EtAl:2016:N16-1}   extends  TransD to additionally use two projection vectors for representing each relation. 
In fact, our STransE  model and TranSparse   \citep{JiLH016} can be viewed as direct extensions of the TransR model, where head and tail entities are associated with their own projection matrices,
rather than using the same matrix for both, as in TransR and CTransR.

Recently, several authors have shown that relation paths between entities in KBs provide richer information and improve the relationship prediction \citep{lin-EtAl:2015:EMNLP1,garciaduran-bordes-usunier:2015:EMNLP,guu-miller-liang:2015:EMNLP,wang-EtAl:2016:P16-13,feng-EtAl:2016:COLING1,Liu:2016:HRW:2911451.2911509,NIPS2016_6098,wei-zhao-liu:2016:EMNLP2016,toutanova-EtAl:2016:P16-1,NguyenCoNLL2016}. In addition,  \newcite{NickelMTG15} reviews other
approaches for learning from KBs and multi-relational data.

\section{Experiments}
For link prediction evaluation, we conduct experiments and compare the
performance of our STransE model with published results on the
benchmark WN18 and FB15k datasets \cite{NIPS2013_5071}. Information
about these datasets is given in Table \ref{tab:datasets}.

\begin{table}[ht]
\centering
\resizebox{8.25cm}{!}{
\begin{tabular}{l|lllll}
\hline
\bf Dataset & \bf \#E & \bf \#R & \bf \#Train & \bf \#Valid & \bf \#Test \\
\hline
WN18 & 40,943 & 18 & 141,442 & 5,000 & 5,000\\
FB15k & 14,951 & 1,345 & 483,142 & 50,000 & 59,071\\
\hline
\end{tabular}
}
\caption{Statistics of the experimental datasets used in this study
  (and previous works). \#E is the number of entities, \#R is the
  number of relation types, and \#Train, \#Valid and \#Test are the
  numbers of triples in the training, validation and test sets,
  respectively. }
\label{tab:datasets}
\end{table}

\begin{table*}[ht]
\centering
\resizebox{16cm}{!}{
\begin{tabular}{l|lll|lll|lll|lll}
\hline
\multirow{3}{*}{\bf Method}& \multicolumn{6}{|c}{\textbf{Raw}}& \multicolumn{6}{|c}{\textbf{Filtered}} \\
\cline{2-13}
& \multicolumn{3}{|c}{\bf WN18} & \multicolumn{3}{|c}{\bf FB15k} & \multicolumn{3}{|c}{\bf WN18} & \multicolumn{3}{|c}{\bf FB15k}\\
\cline{2-13}
\cline{2-13}
&   MR & H10 & MRR &   MR & H10 & MRR &   MR & H10 & MRR &   MR & H10 & MRR \\
\hline
SE \citep{bordes-2011} & 1011 & 68.5 & - & 273& 28.8 & - & 985 &80.5 & - &  162  & 39.8& - \\
Unstructured \citep{Bordes2014SME}  & 315  & 35.3 & - &1074 & 4.5 & - &  304 & 38.2  &-&   979 &  6.3& - \\
 TransE \citep{NIPS2013_5071}  & 263 & 75.4 & - & 243 & 34.9 & - & 251 & 89.2 & -  &  125 & 47.1 & - \\
TransH \citep{AAAI148531}  & 401 & 73.0 & - & 212& 45.7 & -&  303 & 86.7 & - &  87 &  64.4&- \\
TransR \citep{AAAI159571}  & 238& 79.8 & -  & 198&48.2 & -  &  225 &92.0 & - &   77 &   68.7&- \\
CTransR \citep{AAAI159571}  & 231& 79.4 & - & 199& 48.4& - & 218 &  92.3& -  &   75  & 70.2 &  -\\
KG2E \citep{He:2015}  &342& 80.2  & - & \textbf{174} & 48.9& -   &  331 & 92.8 & - &  59 & 74.0& -  \\
TransD \citep{ji-EtAl:2015:ACL-IJCNLP}  & 224 & 79.6 & - & 194 & 53.4 & - &  212 &  92.2 & - &   91 & {77.3}&  -  \\
lppTransD \citep{yoon-EtAl:2016:N16-1} & 283 & 80.5& -  & 195 & 53.0 & - & 270 & 94.3 & - & 78  & 78.7 & -\\
TranSparse \citep{JiLH016} & 223 & 80.1& -  & {187} & \textbf{53.5} & -  & 211  & 93.2 & - & 82  & 79.5& -\\

TATEC \citep{Garcia-DuranBUG15}  & - & -  &- & -&- & -& - &  - & - &   \textbf{58} & 76.7  & -  \\
NTN \citep{NIPS2013_5028}  &- & -  & -& -& -&- & - & {66.1}  & 0.53& - & 41.4& 0.25  \\
DISTMULT \citep{yang-etal-2015}  & -& - & -& -&- &- & - & {{94.2}} & 0.83 & -  & 57.7 & 0.35\\
HolE \citep{Nickel:2016:HEK:3016100.3016172} & -&- &\textbf{0.616} &- &- &0.232& -& \textbf{94.9} & \textbf{0.938} & - & 73.9 & 0.524\\ 
\hline
Our STransE   & \textbf{217} & \textbf{80.9} & 0.469 & 219 & 51.6 & \textbf{0.252} &  \textbf{206}  & {93.4}& 0.657 & 69  & \textbf{79.7}& \textbf{0.543}\\
\hline
\hline
\textsc{r}TransE \citep{garciaduran-bordes-usunier:2015:EMNLP}  & -&  - &- & -&- &- &  -& - & - & {\textbf{50}} & 76.2  & - \\
PTransE \citep{lin-EtAl:2015:EMNLP1}  & -& - & -& 207& 51.4& -  & - & -  & - & {58} & {84.6}& - \\
GAKE \citep{feng-EtAl:2016:COLING1} & - & - & - & 228  &44.5 &-& - & - & - & 119 & 64.8& - \\ 
Gaifman \citep{NIPS2016_6098} & - & - & - & - & - &- & 352  & 93.9 & -& 75 & 84.2 & - \\ 
Hiri \citep{Liu:2016:HRW:2911451.2911509} & - & - & - & - & - & - & -  & 90.8 & 0.691& -  & 70.3& 0.603 \\
\hline
NLFeat \citep{toutanova-chen:2015:CVSC}  &- & - &- & -&- &- & -   & \textbf{94.3} & \textbf{0.940}& -  & \textbf{87.0}& \textbf{0.822} \\
TEKE\_H  \citep{DBLP:conf/ijcai/WangL16} & \textbf{127}  & 80.3 & -& 212  & 51.2 & -&  \textbf{114}  & 92.9 & -& 108 & 73.0& -  \\ 
SSP \citep{0005HZ16} & 168 & \textbf{81.2} & - & \textbf{163} & \textbf{57.2} & - & 156 & 93.2 & - & 82 & 79.0 & -\\ 
\hline
\end{tabular}
}
\caption{Link prediction results. MR, H10 and MRR denote evaluation metrics of mean rank, Hits@10 (in \%) and mean reciprocal rank, respectively. ``NLFeat'' abbreviates Node+LinkFeat. The results for NTN {\protect\citep{NIPS2013_5028}} listed in this table are taken from {\protect\citet{yang-etal-2015}} since  NTN was  originally evaluated on different datasets. }
\label{02tab:linkprediction}
\end{table*}

\subsection{Task and evaluation protocol}
The link prediction task
\cite{bordes-2011,Bordes2014SME,NIPS2013_5071} \CHANGEB{predicts} the head or tail entity given the relation type and
the other \CHANGEB{entity}, i.e. predicting $h$ given $(?, r, t)$ or predicting
$t$ given $(h, r, ?)$ where $?$ denotes the missing element. The results are evaluated using the ranking  induced by the score function $f_r(h,t)$ on test triples.

For each test triple $(h, r, t)$, we corrupted it by replacing either $h$ or $t$ by each of {the possible} 
entities in turn, and then rank these candidates in ascending order of
their implausibility value computed by the score function. This is called as the ``Raw'' setting
protocol. For the ``Filtered'' setting protocol described
in  \citet{NIPS2013_5071}, 
{we  \textit{removed} any corrupted triples}
that appear in the knowledge base,
to avoid cases where a correct corrupted triple might be  ranked higher than the
test triple.  The ``Filtered'' setting thus provides
a clearer view on the ranking performance. 
Following \citet{NIPS2013_5071}, we report the mean rank and the 
Hits@10  (i.e., the proportion of test triples in which
the target entity was ranked in the top 10 predictions) for each model. In addition,  we  report the
mean reciprocal rank, which is commonly
used in information retrieval. 
In both ``Raw'' and
``Filtered'' settings, lower mean
rank, higher mean reciprocal rank or higher Hits@10 indicates better link prediction performance.

Following TransR  \citep{AAAI159571}, TransD
\citep{ji-EtAl:2015:ACL-IJCNLP}, 
\textsc{r}TransE \citep{garciaduran-bordes-usunier:2015:EMNLP},
PTransE \citep{lin-EtAl:2015:EMNLP1}, TATEC \citep{Garcia-DuranBUG15} and TranSparse \citep{JiLH016}, 
we used the entity and relation
vectors produced by TransE \cite{NIPS2013_5071} to initialize the entity
and relation vectors in STransE\CHANGEB{, and} we initialized the relation matrices
with identity matrices.  
We applied the ``Bernoulli'' trick used also in previous work for generating head or tail entities when sampling incorrect triples \citep{AAAI148531,AAAI159571,He:2015,ji-EtAl:2015:ACL-IJCNLP,lin-EtAl:2015:EMNLP1,yoon-EtAl:2016:N16-1,JiLH016}. 
We ran SGD for 2,000 epochs to estimate the
model parameters.  Following \newcite{NIPS2013_5071} we used a grid
search on \CHANGEB{validation set to choose either the $l_1$ or $l_2$ norm in the score function $f$, as well as to} set the SGD learning rate
$\lambda\in\lbrace 0.0001, 0.0005, 0.001, 0.005, 0.01 \rbrace$, the
margin hyper-parameter $\gamma\in\lbrace 1, 3, 5 \rbrace$ and the vector size  $k\in\lbrace 50, 100 \rbrace$.  The lowest filtered mean rank on the
\CHANGEB{validation} set was obtained when \CHANGEB{using the $l_1$ norm in $f$ on both
WN18 and FB15k, and when} $\lambda = 0.0005, \gamma = 5,
\text{ and } k = 50$ for WN18, and $\lambda = 0.0001, \gamma = 1,
\text{ and } k = 100$ for FB15k. 

\subsection{Main results}
Table \ref{02tab:linkprediction} compares the link prediction results of
our STransE model with results reported in prior work, using the same
experimental setup. The first 15 rows report the performance of
the models that do not exploit information about alternative paths between
head and tail entities. The next 5 rows report results of the  models that exploit information about relation paths. The last 3
rows present  results for the models  which make  use of
textual mentions derived from a large external corpus.

It is clear that the models with the additional external corpus
information obtained best results. In future work we plan to extend
the STransE model to incorporate such additional information. Table 
\ref{02tab:linkprediction}  also shows that the models   employing path information generally achieve better results
than models that do not use such information. In terms of models not
exploiting path information or external information, the STransE model
produces the highest filtered mean rank  on WN18 and  the highest
filtered Hits@10 and mean reciprocal rank on FB15k. Compared to the closely related models SE, TransE, TransR,
CTransR, TransD and TranSparse,  our STransE model does better than these models on both WN18
and FB15k.

Following \newcite{NIPS2013_5071}, Table \ref{tab:headtail} analyzes  
Hits@10 results on FB15k with respect to the relation categories
defined as follows: for each relation type $r$, we computed the
averaged number $a_h$ of heads $h$ for a pair $(r, t)$ and the averaged
number $a_t$ of tails $t$ for a pair $(h, r)$. If $a_h < 1.5$ and $a_t
< 1.5$, then $r$ is labeled \textbf{1-1}. If $a_h \geq 1.5$ and $a_t
< 1.5$, then $r$ is labeled \textbf{M-1}. If $a_h < 1.5$ and $a_t
\geq 1.5$, then $r$ is labeled as \textbf{1-M}. If $a_h \geq 1.5$
and $a_t \geq 1.5$, then $r$ is labeled as \textbf{M-M}. 1.4\%,
8.9\%, 14.6\% and 75.1\% of the test triples belong to a relation type
classified as  \textbf{1-1}, \textbf{1-M}, \textbf{M-1} and \textbf{M-M},
respectively.

\begin{table}[ht]
\centering
\resizebox{8cm}{!}{
\setlength{\tabcolsep}{0.25em}
\begin{tabular}{l|llll|llll}
\hline
\multirow{2}{*}{\bf Method}& \multicolumn{4}{|c}{\bf Predicting head $h$} & \multicolumn{4}{|c}{\bf Predicting tail  $t$}\\
\cline{2-9}
& 1-1 & 1-M & M-1 & M-M & 1-1 & 1-M & M-1 & M-M \\
\hline
SE  & 35.6 & 62.6 & 17.2 & 37.5 & 34.9 & 14.6 & 68.3 & 41.3 \\
Unstr. & 34.5 & 2.5 & 6.1 & 6.6  & 34.3 & 4.2 & 1.9 & 6.6\\
 TransE  & 43.7 & 65.7 & 18.2 & 47.2 & 43.7 & 19.7 & 66.7 & 50.0  \\
TransH  & 66.8 & 87.6 & 28.7 & 64.5 & 65.5 & 39.8 & 83.3 & 67.2 \\
TransR  & 78.8 & 89.2 & 34.1 & 69.2 & 79.2 & 37.4 & 90.4 & 72.1 \\
CTransR   & 81.5 & 89.0 & 34.7 & 71.2  & 80.8 & 38.6 & 90.1 & 73.8 \\
KG2E   & \textbf{92.3} & {94.6} & \textbf{66.0} & 69.6 &  \textbf{92.6} & \textbf{67.9} & {94.4} & 73.4  \\
TATEC    & 79.3 & 93.2 & 42.3 & 77.2 & 78.5 & 51.5 & 92.7 & 80.7 \\
TransD   & 86.1 & \textbf{95.5} & 39.8 & 78.5 & 85.4 & 50.6 & {94.4} & 81.2  \\
lppTransD & 86.0 & 94.2 & 54.4 & \textbf{82.2} & 79.7 & 43.2 & \textbf{95.3} & 79.7\\
TranSparse & 86.8 & \textbf{95.5} & 44.3 & {80.9} & 86.6 & 56.6 & 94.4 & \textbf{83.3}\\
\hline
STransE & {82.8} & {94.2} & {50.4} & {80.1} & {82.4} & {56.9} & {93.4} & {83.1}\\ 

\hline
\end{tabular}
}
\caption{Hits@10 (in \%) by the relation category on  FB15k. ``Unstr.'' abbreviates Unstructured.}
\label{tab:headtail}
\end{table}

Table \ref{tab:headtail} shows that in comparison to
prior models not using path information, STransE obtains the second highest
Hits@10 result for \textbf{M-M} relation category at $(80.1\% + 83.1\%) / 2 = 81.6\%$  which is  0.5\% smaller than the Hits@10 result of TranSparse  for \textbf{M-M}. However, STransE obtains 2.5\% higher Hits@10 result than TranSparse for \textbf{M-1}. In
addition, STransE also performs better than TransD for \textbf{1-M} and
\textbf{M-1} relation categories. We believe the improved performance of
the STransE model is due to its use of full matrices\CHANGEB{,} rather than \CHANGEB{just}
projection vectors as in TransD. This permits STransE to model diverse
and complex relation categories (such as \textbf{1-M}, \textbf{M-1} and
especially \textbf{M-M}) better than TransD \CHANGEA{and other \CHANGEB{similiar} models}. However, STransE is not
as good as TransD for the \textbf{1-1} \CHANGEB{relations}. 
\CHANGEB{Perhaps the extra parameters in STransE hurt performance in this case
(note that 1-1 relations are relatively rare, so STransE does better overall).}

\section{Conclusion and future work}


This paper presented a new embedding model for link prediction and KB completion. \CHANGEA{Our STransE combines insights \CHANGEB{from} several simpler embedding models, specifically \CHANGEB{the} Structured Embedding \CHANGEB{model} \cite{bordes-2011} and \CHANGEB{the} TransE \CHANGEB{model} \cite{NIPS2013_5071}}\CHANGEB{,} by using a low-dimensional vector and two \CHANGEB{projection} matrices to represent each relation. STransE, \CHANGEA{while being conceptually simple}, produces highly competitive results on standard link prediction evaluations, \CHANGEA{and scores better than the embedding-based models it builds on. Thus it is a suitable candidate for serving as future baseline for more complex models in the link prediction task.}

 In future work we plan to extend STransE to exploit relation path information in knowledge bases, in a manner similar to \newcite{lin-EtAl:2015:EMNLP1}, 
\newcite{guu-miller-liang:2015:EMNLP} or \newcite{NguyenCoNLL2016}.

\section*{Acknowledgments}  
This research was supported by a Google award through the Natural 
Language Understanding Focused Program, and under the Australian 
Research Council's {\em Discovery Projects} funding scheme (project 
number DP160102156).

NICTA is funded by the Australian Government through the Department of Communications and the Australian Research Council through the ICT Centre of  Excellence Program. The first author is supported by an International Postgraduate Research Scholarship and a NICTA NRPA Top-Up Scholarship. 

\bibliographystyle{naaclhlt2016}
\bibliography{REFs}

\end{document}